# QERA: an Analytical Framework for Quantization Error Reconstruction

### v1

@@ IV-B Improved QPEFT
TABLE I: Fine-tuning results of RoBERTa-base on GLUE. QERA-approx outperforms LoftQ across all bit widths, and the improvement is more obvious with aggressive quantization. QERA achieves $\Delta_{\text{acc}}$ = 4.12% higher than LoftQ at 3-bit and 6.05% at 2-bit.
@@
TABLE II: Fine-tuning results of LLaMA-2-7B and LLaMA-3.1-8B on SlimPajama and GSM8K. A trend similar to RoBERTa experiments are observed, i.e., QERA outperforms QLoRA and LoftQ and the improvement is more obvious on aggressive quantization.
@@
Figure 2: Faster convergence of QERA-approx on STSB.
@@
This empirical evidence suggests a strong correlation between the reduction of layer output error and the decrease in model output error. Conversely, minimizing weight approximation error does not demonstrate a comparable impact on overall model performance.
@@
Table I and Table II summarize the fine-tuning experiments of RoBERTa-base on GLUE, and LLaMA-2-7B/-3.1-8B fine-tuned on SlimPajama and GSM8K, respectively. QERA outperforms both LoftQ and QLoRA. In GLUE experiments, at 4-bit, QERA enables an average accuracy gain of 0.96% and 0.79% higher than QLoRA and LoftQ respectively, close to BF16 LoRA; At 3-bit and 2-bit, QERA achieves a 4.12% and 6.05% higher average accuracy than LoftQ respectively. Similar trends are observed on LLM fine-tuning experiments, i.e., QERA outperforms QLoRA and LoftQ, and the advantage of QERA over LoftQ is more obvious with more aggressive quantization.
@@ IV-C Improved PTQ
As mentioned at the end of Section III-C, the scale matrix in LQER [4] is similar to the one in QERA-approx, but is based on hand-crafted heuristics. As a result, we observe that the model performance of LQER varies randomly as the number of calibration samples increases (the purple curve in Figure 3). On the contrary, more calibration samples consistently lead to better model performance for QERA until convergence.
Improved perplexity and downsteam task accuracy
We apply QERA-approx and QERA-exact to a range of models and evaluate on both pretraining task and downstream tasks in Table III and Table IV respectively. We also compare to HQQ, a SoTA method that does not use quantization error reconstruction.
TABLE III: Perplexity ($\downarrow$) of LLMs on WikiText2. $w$-only denotes the quantized model without low-rank error reconstruction. QERA-approx outperforms LQER on almost all setups and QERA-exact achieves the lowest perplexity. The advantage of QERA is pronounced at 3-bit.
@@
TABLE IV: Average accuracy ($\uparrow$) of LLMs on six downstream tasks. QERA-exact outperforms other quantization-error reconstruction-based methods across almost all models. We also compare to HQQ [23], a SoTA PTQ method that does not adopt quantization-error reconstruction or activation heuristics. QERA-exact achieves an average accuracy on par with HQQ.
@@
On most models, QERA-approx outperforms ZeroQuant-V2 and LQER, while QERA-exact achieves the best performance. At 4-bit, QERA-exact is nearly lossless. At 3-bit, QERA-exact’s improvement over QERA-approx (Table III) is clear, indicating the superiority of QERA-exact for aggressive quantization.
@@
To better understand the impact on instruction-tuned models, we present the results of Vicuna-7b-v1.5 on AlpacaEval 2.0. In Figure 4, we evaluate the quantization-error-reconstruction-based methods against the $w$-only quantization counterpart. QERA outperforms ZeroQuant-V2 and LQER by a higher win rate, indicating a better response quality.
V Discussion
In this section, we revisit the arguments, design choices, and observations made in the previous sections, including a test of Assumption 1, and the choice of the calibration set for PEFT. An extended discussion of numeric stability and scalability of QERA is in Section A-G, and LoRA rank and model choices of PEFT experiments is in Section A-H.
@@ V Discussion
To test Assumption 1, we profile the autocorrelation matrix ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}$ of the linear layer inputs in LLaMA-2-7B and LLaMA-3-8B. Note that ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}_{i,j}=\mathbb{E}_{{\bm{x}}\sim{\mathbb{X}}}\{x_{i}x_{j}\}$, which assumes to be zero for $i\neq j$ in Assumption 1.  Figure 5 shows the normalized ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}$ magnitude, $\frac{\mathrm{abs}(\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}})}{||\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}||_{F}}$, of four representative layers in LLaMA-3-8B where darker elements denote values closer to zero. There are several layers with some input dimensions strongly correlated with others, such as the inputs to the third attention layer in Figure 5(a), but for most layers, our assumption holds, especially the MLP layers, such as Figures 5(b), 5(c) and 5(d). More ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}$ plots are in Section A-E.
Choice of calibration set for QPEFT
One problem is to determine the calibration set for QERA before fine-tuning. In 2-bit RoBERTa-base fine-tuning experiment on SST2 (Section A-F), we find that calibrating on the pretraining dataset, WikiText2, helps the loss to decrease. However, the loss of the model calibrated on the fine-tuning dataset does not follow the same trend. We hypothesize that the massive padding tokens in preprocessed SST2 samples cause this discrepancy, especially considering that the sequence length of the raw SST2 dataset changes fiercely.
@@
VI Conclusion
@@ Appendix A Appendix
A-A Algorithms in Related Work
Here we summarize the algorithm of LoftQ [1] in Algorithm 1 and LQER [4] in Algorithm 2 respectively. LQ-LoRA [3] adopts a variant of Algorithm 1. ZeroQuant-V2 [2] can be considered as Algorithm 1 with one iteration, or a special case of Algorithm 2 where the scale matrix ${\bm{S}}$ is an identity matrix.
Algorithm 1 LoftQ [1]
0:  Pretrained weight ${\bm{W}}$, target rank $k$, quantization function $\mathrm{q}(\cdot)$, dequantization function $\mathrm{dq}(\cdot)$, number of iterations $T$
1:  ${\bm{A}}_{k}\leftarrow\mathbf{0},{\bm{B}}_{k}\leftarrow\mathbf{0}$
2:  for $i$ = $1$ to $T$ do
3:     ${\bm{W}}_{q}\leftarrow\mathrm{q}({\bm{W}}-{\bm{A}}_{k}{\bm{B}}_{k})$Update quantized weight matrix}
4:     $\widetilde{{\bm{W}}}\leftarrow\mathrm{dq}({\bm{W}}_{q})$
5:     ${\bm{U}},{\bm{\Sigma}},{\bm{V}}^{T}\leftarrow\mathrm{SVD}({\bm{W}}-\widetilde{{\bm{W}}})$SVD-based rank-$k$ approximation}
6:     ${\bm{A}}_{k}\leftarrow{\bm{U}}_{:,:k}\sqrt{{\bm{\Sigma}}_{:k,:k}}$, ${\bm{B}}_{k}\leftarrow\sqrt{{\bm{\Sigma}}_{:k,:k}}{\bm{V}}^{T}_{:k,:}$
7:  end for
Algorithm 2 LQER [4]
0:  Pretrained weight ${\bm{W}}$, target rank $k$, quantization function $\mathrm{q}(\cdot)$, dequantization function $\mathrm{dq}(\cdot)$, calibration dataset ${\mathbb{X}}=\{{\bm{x}}_{i}\in{\mathbb{R}}^{m}|i=1,\dots,N\}$
1:  Initialize vector ${\bm{s}}\leftarrow\mathbf{0}$
2:  for sample ${\bm{x}}$ in ${\mathbb{X}}$ do
3:     ${\bm{s}}\leftarrow{\bm{s}}+\mathrm{abs}({\bm{x}})$Calibrate by accumulating activation magnitude on each dimension}
4:  end for
5:  ${\bm{S}}\leftarrow\frac{1}{N}\mathrm{diag}({\bm{s}})$Construct a diagonal matrix ${\bm{S}}$}
6:  ${\bm{W}}_{q}\leftarrow\mathrm{q}({\bm{W}})$
7:  $\widetilde{{\bm{W}}}\leftarrow\mathrm{dq}({\bm{W}}_{q})$
8:  ${\bm{U}},{\bm{\Sigma}},{\bm{V}}^{T}\leftarrow\mathrm{SVD}({\bm{S}}({\bm{W}}-\ \widetilde{{\bm{W}}}))$SVD on the scaled weight error}
9:  ${\bm{A}}_{k}\leftarrow{\bm{S}}^{-1}{\bm{U}}_{:,:k}$, ${\bm{B}}_{k}\leftarrow{\bm{\Sigma}}_{:k,:k}{\bm{V}}^{T}_{:k,:}$Rank-$k$ approximation with un-scaling}
A-B Proof of Theorem 2
@@
A-C Detailed Experiment Setup
We mainly use PyTorch, Transformers, PEFT, and Accelerate to implement QERA. We use SciPy’s implementation of blocked Schur algorithm [36] to calculate the matrix square root, which runs on CPUs. The evaluation is performed with lm-evaluation-harness, Evaluate, and AlpacaEval 2.0 [35].
A-C1 QPEFT Hyperparameters

### v2

@@ 4.2 Improved QPEFT
Table 1: Fine-tuning results of RoBERTa-base on GLUE. QERA-approx outperforms LoftQ across all bit widths, and the improvement is more obvious with aggressive quantization. QERA achieves $\Delta_{\text{acc}}$ = 4.12% higher than LoftQ at 3-bit and 6.05% at 2-bit.
@@
Table 2: Fine-tuning results of LLaMA-2-7B and LLaMA-3.1-8B on SlimPajama and GSM8K. A trend similar to RoBERTa experiments are observed, i.e., QERA outperforms QLoRA and LoftQ and the improvement is more obvious on aggressive quantization.
@@
This empirical evidence suggests a strong correlation between the reduction of layer output error and the decrease in model output error in QER problem. Conversely, minimizing weight approximation error using LoftQ does not have a comparable impact on overall model performance.
Figure 2: Faster convergence of QERA-approx on STSB.
@@
Table 1 and Table 2 summarize the fine-tuning experiments of RoBERTa-base on GLUE, and LLaMA-2-7B/-3.1-8B fine-tuned on SlimPajama and GSM8K, respectively. QERA outperforms both LoftQ and QLoRA. In GLUE experiments, at 4-bit, QERA enables an average accuracy gain of 0.96% and 0.79% higher than QLoRA and LoftQ respectively, close to BF16 LoRA; At 3-bit and 2-bit, QERA achieves a 4.12% and 6.05% higher average accuracy than LoftQ respectively. Similar trends are observed on LLM fine-tuning experiments, i.e., QERA outperforms QLoRA and LoftQ, and the advantage of QERA over LoftQ is more obvious with more aggressive quantization.
@@ 4.3 Improved PTQ
As mentioned at the end of Section 3.3, the scale matrix in LQER (Zhang et al., 2024a) is similar to the one in QERA-approx, but is based on hand-crafted heuristics. As a result, we observe that the model performance of LQER varies randomly as the number of calibration samples increases (the purple curve in Figure 3). On the contrary, more calibration samples consistently lead to better model performance for QERA until convergence.
Table 3: Perplexity ($\downarrow$) of LLMs on WikiText2. $w$-only denotes the quantized model without low-rank error reconstruction. QERA-approx outperforms LQER on almost all setups and QERA-exact achieves the lowest perplexity. The advantage of QERA is pronounced at 3-bit.
@@
Improved perplexity and downsteam task accuracy
We apply QERA-approx and QERA-exact to a range of models and evaluate on both pretraining task and downstream tasks in Table 3 and Table 4 respectively. We also compare to HQQ, a SoTA method that does not use quantization error reconstruction. On most models, QERA-approx outperforms ZeroQuant-V2 and LQER, while QERA-exact achieves the best performance. At 4-bit, QERA-exact is nearly lossless. At 3-bit, QERA-exact’s improvement over QERA-approx (Table 3) is clear, indicating the superiority of QERA-exact for aggressive quantization.
Table 4: Average accuracy ($\uparrow$) of LLMs on six downstream tasks. QERA-exact outperforms other quantization-error reconstruction-based methods across almost all models. We also compare to HQQ (Badri & Shaji, 2023), a SoTA PTQ method that does not adopt quantization-error reconstruction or activation heuristics. QERA-exact achieves an average accuracy on par with HQQ.
@@
To better understand the impact on instruction-tuned models, we present the results of Vicuna-7b-v1.5 on AlpacaEval 2.0. In Figure 4, we evaluate the QER-based methods against the $w$-only quantization counterpart. QERA outperforms ZeroQuant-V2 and LQER by a higher win rate, indicating a better response quality.
5 Discussion
In this section, we revisit the arguments, design choices, and observations made in the previous sections, including a test of Assumption 1, and the choice of the calibration set for PEFT. We offer an extended discussion of the numeric stability and scalability in Section A.7, and LoRA rank and model choices of PEFT experiments in Section A.9.
@@ 5 Discussion
To test Assumption 1, we profile the autocorrelation matrix ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}$ of the linear layer inputs in LLaMA-2-7B and LLaMA-3-8B. Note that ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}_{i,j}=\mathbb{E}_{{\bm{x}}\sim{\mathbb{X}}}\{x_{i}x_{j}\}$, which assumes to be zero for $i\neq j$ in Assumption 1.  Figure 5 shows the normalized ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}$ magnitude, $\frac{\mathrm{abs}(\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}})}{||\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}||_{F}}$, of four representative layers in LLaMA-3-8B where darker elements denote values closer to zero. There are several layers with some input dimensions strongly correlated with others, such as the inputs to the third attention layer in Figure 5(a), but for most layers, our assumption holds, especially the MLP layers, such as Figures 5(b), 5(c) and 5(d). More ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}$ plots are in Section A.11.
@@
Choice of calibration set for QPEFT
One problem is to determine the calibration set for QERA before fine-tuning. In 2-bit RoBERTa-base fine-tuning experiment on SST2 (Section A.6), we find that calibrating on the pretraining dataset, WikiText2, helps the loss to decrease. However, the loss of the model calibrated on the fine-tuning dataset does not follow the same trend. We hypothesize that the massive padding tokens in preprocessed SST2 samples cause this discrepancy, especially considering that the sequence length of the raw SST2 dataset changes fiercely.
6 Conclusion
@@
Acknowledgments
This work was sponsored by Advanced Research + Invention Agency (ARIA), UK. We also thank ARIA for their research network.
@@ Appendix A Appendix
A.1 Algorithms in Related Work
Here we summarize the algorithm of LoftQ (Li et al., 2023) in Algorithm 1 and LQER (Zhang et al., 2024a) in Algorithm 2 respectively. LQ-LoRA (Guo et al., 2023) adopts a variant of Algorithm 1. ZeroQuant-V2 (Yao et al., 2023) can be considered as Algorithm 1 with one iteration, or a special case of Algorithm 2 where the scale matrix ${\bm{S}}$ is an identity matrix.
Algorithm 1 LoftQ (Li et al., 2023)
1:Pretrained weight ${\bm{W}}$, target rank $k$, quantization function $\mathrm{q}(\cdot)$, dequantization function $\mathrm{dq}(\cdot)$, number of iterations $T$
2:${\bm{A}}_{k}\leftarrow\mathbf{0},{\bm{B}}_{k}\leftarrow\mathbf{0}$
3:for $i$ = $1$ to $T$ do
4:     ${\bm{W}}_{q}\leftarrow\mathrm{q}({\bm{W}}-{\bm{A}}_{k}{\bm{B}}_{k})$triangleright Update quantized weight matrix
5:     $\widetilde{{\bm{W}}}\leftarrow\mathrm{dq}({\bm{W}}_{q})$
6:     ${\bm{U}},{\bm{\Sigma}},{\bm{V}}^{T}\leftarrow\mathrm{SVD}({\bm{W}}-\widetilde{{\bm{W}}})$triangleright SVD-based rank-$k$ approximation
7:     ${\bm{A}}_{k}\leftarrow{\bm{U}}_{:,:k}\sqrt{{\bm{\Sigma}}_{:k,:k}}$, ${\bm{B}}_{k}\leftarrow\sqrt{{\bm{\Sigma}}_{:k,:k}}{\bm{V}}^{T}_{:k,:}$
8:end for
Algorithm 2 LQER (Zhang et al., 2024a)
1:Pretrained weight ${\bm{W}}$, target rank $k$, quantization function $\mathrm{q}(\cdot)$, dequantization function $\mathrm{dq}(\cdot)$, calibration dataset ${\mathbb{X}}=\{{\bm{x}}_{i}\in{\mathbb{R}}^{m}|i=1,\dots,N\}$
2:Initialize vector ${\bm{s}}\leftarrow\mathbf{0}$
3:for sample ${\bm{x}}$ in ${\mathbb{X}}$ do $\triangleright$ Calibration
4:     ${\bm{s}}\leftarrow{\bm{s}}+\mathrm{abs}({\bm{x}})$triangleright Accumulate activation magnitude on each dimension
5:end for
6:${\bm{S}}\leftarrow\frac{1}{N}\mathrm{diag}({\bm{s}})$triangleright Construct a diagonal matrix ${\bm{S}}$
7:${\bm{W}}_{q}\leftarrow\mathrm{q}({\bm{W}})$
[…]
9:${\bm{U}},{\bm{\Sigma}},{\bm{V}}^{T}\leftarrow\mathrm{SVD}({\bm{S}}({\bm{W}}-\ \widetilde{{\bm{W}}}))$triangleright SVD on the scaled weight error
10:${\bm{A}}_{k}\leftarrow{\bm{S}}^{-1}{\bm{U}}_{:,:k}$, ${\bm{B}}_{k}\leftarrow{\bm{\Sigma}}_{:k,:k}{\bm{V}}^{T}_{:k,:}$triangleright Rank-$k$ approximation with un-scaling
A.2 Proof of Theorem 2
@@
A.3 Connection and Difference between CALDERA and QERA
CALDERA (Saha et al., 2024) is the concurrent work close to QERA. Here we elaborate the connection and difference between CALDERA and QERA, and highlight the contributions of QERA.
CALDERA focuses on a different problem setup. Specifically, CALDERA focuses on the following problem:
$\mathrm{min}_{\widetilde{{\bm{W}}},{\bm{A}}_{k,q},{\bm{B}}_{k,q}}||{\bm{X}}{\bm{W}}-{\bm{X}}(\widetilde{{\bm{W}}}+{\bm{A}}_{q,k}{\bm{B}}_{q,k})||_{F}^{2}$ (32)
where ${\bm{X}}\in{\mathbb{R}}^{b\times m}$ denotes a batch of calibration samples, and $\widetilde{{\bm{W}}}$, ${\bm{A}}_{k,q}$, and ${\bm{B}}_{k,q}$ are all quantized variables to optimize. Note that this problem setup is different from QERA (Equation 9):
where only the low-rank high-precision ${\bm{C}}_{k}:={\bm{A}}_{k}{\bm{B}}_{k}$ is the variable to optimize, and the quantized weight $\widetilde{{\bm{W}}}$ is predefined given a quantization method.
Table 5: Notation Table for the Equivalence Derivation
Notation Description Comments
$b$ Number of calibration samples (vectors)
$m$ Layer input feature size
$n$ Layer output feature size
${\bm{X}}$ Calibration set Shape: $b\times m$
${\bm{x}}$ A sample in the calibration set Shape: $1\times m$
${\bm{W}}$ Original full-precision layer weights Shape: $m\times n$
${\bm{Y}}$ Layer output matrix corresponding to ${\bm{X}}$ Shape: $b\times n$
${\bm{y}}$ Layer output vector corresponding to ${\bm{x}}$ Shape: $1\times n$
$k$ Rank of the low-rank approximation
${\bm{C}}_{k}$ Approximated rank-$k$ weight Shape: $m\times n$
${\bm{U}},{\bm{\Sigma}},{\bm{V}}$ SVD decomposition of ${\bm{X}}$
$\mathrm{SVD}_{k}(\cdot)$ Truncated rank-$k$ SVD
We find that CALDERA’s Lemma 4.2 is equivalent to Theorem 1 in QERA. Note that the proof of QERA-exact is different from Caldera’s Lemma 4.2, though the final closed-form solution is equivalent. Here we additionally show the derivation of the equivalence between QERA-exact and Caldera’s Lemma 4.2 using the notation table in Table 5. For convenience, we remove the quantized weight term $\widetilde{{\bm{W}}}$ from QERA (Problem 2 in Equation 9), which does not change the proof. Now the problem becomes finding the optimal low-rank approximation of the weight matrix, ${\bm{C}}_{k}$ that minimizes the layer output error.
First we note that the objective of QERA, Equation 9, is equivalent to CALDERA’s Eq(5) scaled by a constant $n$:
$\begin{split}&\text{QERA}:\text{min}_{{\bm{C}}_{k}}E_{\mathbf{x}}\{||\mathbf{x}({\bm{C}}_{k}-{\bm{W}})||_{2}^{2}\}\\ &\text{CALDERA}:\text{min}_{{\bm{C}}_{k}}||{\bm{X}}({\bm{C}}_{k}-{\bm{W}})||_{F}^{2}\end{split}$ (34)
Then we show that Theorem 1 (QERA-exact) is equal to Caldera’s Lemma 4.2.
[…]
$\text{CALDERA}:{\bm{C}}_{k}^{\prime}={\bm{V}}{\bm{\Sigma}}\cdot\mathrm{SVD}_{k}({\bm{U}}^{T}{\bm{Y}})$ (36)
We firstly show that $({\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}})^{-1}$ in Equation 35 equals to ${\bm{V}}{\bm{\Sigma}}$ in Equation 36 scaled by a constant ${\sqrt{b}}$:
$\begin{split}{\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}&=\frac{1}{b}({\bm{X}}^{T}{\bm{X}})={\bm{V}}{\bm{\Sigma}}{\bm{U}}^{T}{\bm{U}}{\bm{\Sigma}}{\bm{V}}^{T}={\bm{V}}{\bm{\Sigma}}^{2}{\bm{V}}^{T}\\ {\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}&=\frac{1}{\sqrt{b}}{\bm{\Sigma}}{\bm{V}}^{T}\\ ({\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}})^{-1}&=\sqrt{b}{\bm{V}}{\bm{\Sigma}}^{-1}\end{split}$ (37)
Then we show that ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}{\bm{W}}$ in Equation 35 equals to ${\bm{U}}^{T}{\bm{Y}}$ in Equation 36 scaled by the constant $\frac{1}{\sqrt{b}}$:
$\begin{split}{\bm{U}}^{T}{\bm{Y}}&={\bm{U}}^{T}{\bm{X}}{\bm{W}}={\bm{U}}^{T}{\bm{U}}{\bm{\Sigma}}{\bm{V}}^{T}{\bm{W}}={\bm{\Sigma}}{\bm{V}}^{T}{\bm{W}}={\sqrt{b}}{\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}{\bm{W}}\\ {\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}^{\frac{1}{2}}{\bm{W}}&=\frac{1}{\sqrt{b}}{\bm{U}}^{T}{\bm{Y}}\end{split}$ (38)
Therefore $C_{k}$ equals to $C_{k}^{\prime}$, and the two solutions are equivalent. Despite of the equivalence, we shortlist the differences between CALDERA and our work:
  • Different problem setup (Equation 34).
  • We simplify QERA-exact and derive QERA-approx, which is a computationally-efficient approximated solution. Specifically, QERA-approx is more suitable for parameter-efficient fine-tuning than QERA-exact/CALDERA. Moreover, QERA-approx overcomes the pitfalls in existing methods and explains why previous heuristic methods like LQER work.
  • The optimization objective is similar (vector form vs matrix form), and the final closed-form solution is equivalent, but the proof of QERA-exact is different from CALDERA.
A.4 Detailed Experiment Setup
We mainly use PyTorch, Transformers, PEFT, and Accelerate to implement QERA. We use SciPy’s implementation of blocked Schur algorithm (Deadman et al., 2012) to calculate the matrix square root, which runs on CPUs. The evaluation is performed with lm-evaluation-harness, Evaluate, and AlpacaEval 2.0 (Dubois et al., 2024).
A.4.1 QPEFT Hyperparameters